A car moves distance of 200 m. It covers first half of the distance at speed 60 kmh–1 and the second half at the speed v. If the average speed is 40 kmh–1, the value of v is

Grade:11

1 Answers

Deepanshu kajla
2 years ago
Total distance =200m=0.2km
Avg speed =total distance /total time 
                     =0.2/0.1/60+0.1/40
                      =0.2*60*40/4+6=24*2=48km/hr